OpenStudy (anonymous):

((b-1) (-axb)^(b-2))/(-bax)^(b-1)

OpenStudy (ranga):

Is this the expression?\[\Large \frac{ (b-1)(-axb)^{b-2} }{ (-bax)^{b-1} }\]

OpenStudy (anonymous):

yes

OpenStudy (ranga):

You can use the identity: x^m / x^n = x^(m-n)

OpenStudy (anonymous):

ok cool, thanks

OpenStudy (ranga):

you are welcome.

OpenStudy (ranga):

= (b-1) * (-axb)^{ (b-2) - (b - 1) } = (b-1) * (-axb)^(-1) = (b-1) / (-axb) = (1-b) / axb